Calculate the percentage yield of this reaction of 2MgO(s) if 0,15 moles of magnesium react with excess oxygen gas to form 5,4 grams of Magnesium oxide
1
Expert's answer
2020-06-29T07:48:08-0400

Given Moles of Mgo is 0.15 moles .

Mg react with O2 and form MgO .


2 mole of Mg gives 2 mole of Mgo .


So 0.15 moles of Mg Gives 0.15 moles of MgO


So , Mass Of MgO = 0.15 * 40 = 6 gm.


So , Maximum Mass of MgO = 6 gm . (Theoretical Mass ).


Experimental Mass of MgO = 5.4 gm .


% Yield (percentage yield) = ( Experimental Mass / Theoretical Mass ) * 100 .

 

percentage yield = (5.4 / 6) * 100 = 90 % .